# Official-NV: An LLM-Generated News Video Dataset for Multimodal Fake News Detection

### v2

@@ 3 CONCLUSION
We propose a video dataset for MFND, constructed using LLM data augmentation. Following this, we furnish a comprehensive description of the creation process of Official-NV and delve into the specifics of the data entities’ attributes and distribution. To assess the dataset’s efficacy, we carry out numerical experiments, evaluating the performance of various baselines and exploring the significance of each modality. Upon comparing the results, it becomes evident that the baselines exhibit strong performance on Official-NV, with the text modality proving to be particularly influential. Looking ahead, we intend to explore additional avenues that hold promise in enhancing the predictive power of this dataset, with the aspiration that it will pave the way for further advancements in MFND research.

### v3

@@ V CONCLUSION
In this paper, we propose a video dataset for MFND, constructed using LLM data augmentation. Following this, we furnish a comprehensive description of the creation process of Official-NV and delve into the specifics of the data entities’ attributes and distribution. To evaluate the effectiveness of the dataset, we introduce a new baseline model, OFNVD, and conduct numerical experiments to assess the performance of various baselines, demonstrating the superiority of Official-NV. In future, we intend to explore additional avenues that hold promise for enhancing the predictive power of this dataset, with the aspiration that it will pave the way for further advancements in MFND research.